# Circuit Component Reuse Across Tasks in Transformer Language Models

### v2

@@ 7.1 Conclusion
A major concern with the methods of mechanistic interpretability we address in this work is the possibility that analyzing circuits in smaller models and/or on toy tasks might not lead to better understanding of neural networks in general. Our work provides a proof of concept where results from an analysis done on a small model (GPT2-Small) on a toy task (IOI) in Wang et al. (2022) allows us to make accurate predictions about a different and more complex task task on a bigger version of that model (GPT2-Medium). Aligning the IOI circuit onto GPT2-Medium revealed a simple and obvious avenue via which we could ‘ideally’ intervene on the model to fix model errors. We believe this result provides a promising avenue for future circuit analysis. Although the scale of the model tested here is still relatively small, related work has already shown that circuit analysis techniques can help us understand models in the billions of parameters range (Lieberum et al., 2023). That, combined with the insights we present here, suggests an exciting path forward for understanding, controlling, and improving even production-scale models.
@@ 7.2 Predicting Model Behaviors
The behavior of LMs is infamously hard to predict. If we know what the core algorithmic steps that LMs employ are, and know about how they tend to decompose complex functions into those steps, we can better anticipate their behavior, and predict what problems they will and won’t solve well, as well as what mistakes they will make. Current auditing work only catches ‘bad’ behavior after a model has already been deployed. A core promise of interpretability research is the ability to understand LMs at a high enough level that we can predict how changes will make them behave. Our work shows an interesting case where a large part of two seemingly different tasks are solved with the same process, and supports manual interventions based on a higher level understanding of model behavior built on lower level analyses.
Connecting similarities in low-level analyses like circuits can lead us to understanding more abstract processes that underlie model behaviors, like those described in Holtzman et al. (2023). For example, knowledge about general interactions between inhibition and mover heads could dispel some of the mystery around model predictions in any task that require copying a context token from multiple possible options. An important future direction in this work could focus on quantifying some measure of circuit similarity, which remains difficult without a better understanding of what we determine as sufficiently explaining some behavior.
@@ A.2 Path Patching
Appendix B Path Patching Results on IOI
Figure 6: Full path patching results for the IOI task. Arrows show the three path patching iterations. Color bars show percent logit difference. The lower the number, the more important the head is, as patching it causes the logit difference to drop (i.e., harms performance). First the mover heads influence the logits. Path patching to the mover head query vectors (to see how these heads ‘decide’ where to attend) reveals the inhibition heads (12.3, 13.4, 13.14). Path patching to the value vectors of the inhibition heads (to see what the inhibition heads write into the residual stream), we find the induction and duplicate token heads, which are sensitive to the duplication of the subject token.
In this section we describe the process we use to reverse-engineer the IOI task on GPT2-Medium. We follow the path patching process from Wang et al. (2022) as closely as possible and are able to replicate the same circuit.
B.1 Mover Heads
@@
Figure 7: Mover head attention patterns showing copying behavior towards the IO token. The negative mover head 19.1 is shown at the bottom with the opposite trend: attending to the subject token and demoting it.
B.2 The Negative Mover Head
@@ B.3 Inhibition Heads
Appendix C Path Patching Results on Colored Objects
@@ C.1 Mover Heads
We show examples of mover heads on the Colored Objects task in Figure 11, which shows that the dot product between the token embedding for color words being attended to and the output of the mover head is high, and correlates positively with the attention paid to color words. That is: these heads copy whatever they attend to, and in this setting, they attend to the color words in the test example. Notice that these are the same heads as in IOI (see §B.1).
@@
[…]
@@
C.2 Content Gatherer Heads
[…]
C.3 What are the Inhibition Heads Doing?
@@
[…]
C.4 What is the Negative Mover Head doing?
@@
Appendix D Duplicate Token and Induction Heads
@@
Appendix E Additional Sources of Error in Colored Objects
@@
E.1 Bias Towards Position in a Sequence
@@
Appendix F Additional Details on Intervention Results
@@
Appendix G Details on Colored Objects Dataset Construction
@@
Appendix H Circuit Overlap in Additional Tasks
@@
H.1 Greater Than Circuit
@@
H.2 World Capitals Circuit:
@@
H.3 Challenges in Quantifying Circuit Overlap
@@
Appendix I Do Other Models also Use Similar Circuits for These Tasks?
@@
Appendix J How was the Colored Objects Task Selected?

### v3

@@ 7.1 Conclusion
A major concern with the methods of mechanistic interpretability we address in this work is the possibility that analyzing circuits in smaller models and/or on toy tasks might not lead to better understanding of neural networks in general. Our work provides a proof of concept where results from an analysis done on a small model (GPT2-Small) on a toy task (IOI) in Wang et al. (2022) allows us to make accurate predictions about a different and more complex task task on a bigger version of that model (GPT2-Medium). Aligning the IOI circuit onto GPT2-Medium revealed a simple and obvious avenue via which we could ‘ideally’ intervene on the model to fix model errors. Although the scale of the model tested here is still relatively small, related work has already shown that circuit analysis techniques can help us understand models in the billions of parameters range (Lieberum et al., 2023). That, combined with the insights we present here, suggests an exciting path forward for understanding, controlling, and improving even production-scale models.
@@ 7.2 Predicting Model Behaviors
The behavior of LMs is infamously hard to predict. Current auditing work only catches ‘bad’ behavior after a model has already been deployed. If we know what the core algorithmic steps that LMs employ are and how they tend to decompose complex functions into those steps, we can better anticipate their behavior and predict what problems they will and won’t solve well. A core promise of interpretability research is the ability to understand LMs at a high enough level that we can predict how changes will make them behave on new data. Our work shows an interesting case where we achieve a higher level understanding of a larger model behavior built on lower level analyses of a small model, though Appendix J suggests that showing this might not be trivial.
Connecting similarities in low-level analyses like circuits can lead us to understanding more abstract processes that underlie model behaviors, like those described in Holtzman et al. (2023). For example, knowledge general inhibition-mover interactions could help explain selection among multiple alternatives in a context.
@@ A.2 Path Patching
Appendix B Glossary of Terms
Throughout the paper, we use terms that are highly specific to this line of work. Here we take more space to provide definitions of such terms.
  1. 1.
    Circuit Although specifics differ depending on the context of its use, a circuit typically refers to the path through the components (attention heads, MLPs) of a model that compute some behavior. If we think of all of the connections through a neural network as edges in a graph, a circuit is the path through those edges that allows the model to solve some task. Olah et al. (2020) define a circuit in much more detail, although the definition used here (and more generally in other work) is a bit looser.
  2. 2.
    ”residual stream” / ”write in the residual stream” Transformers consist of attention and MLP blocks (ignoring LayerNorms). The outputs of these blocks are added to the inputs of the blocks through the residual connection. Since the next word prediction starts as the embedding for the current word, all updates to it are made through residual connections. Therefore, a useful abstraction for how information moves is through a residual stream (Elhage et al., 2021) where information is added to the stream by attention/MLP layers. Therefore writing to the residual stream is done by adding a vector to it. This perspective helps in understanding how path patching is implemented.
  3. 3.
    [end] refers to the last token in a sequence. For example, the “ to” token in the IOI examples.
  4. 4.
    patching to the X vectors” refers to path patching the outputs of earlier heads from one input text into the forward pass of a minimally different input text such that it affects the X vectors of some later head, where X is either the keys, queries, or values of the later head. In other words, we only allow the output of some earlier head to affect the queries/keys/values of the later head.
  5. 5.
    Dot w/ Token Embed (E.g., Figure 4). This measurement is short for ”dot product with the embedding vector of a given token”. In Figures 4 and 11, for example this refers to taking the dot product between the output of some attention head and the vector in the unembedding matrix for some token t of interest. This tells us how much a specific head writes in the direction of a certain token, thus promoting it in the output space. See Wang et al. (2022) for more details.
  6. 6.
    Attn Prob on Token (E.g., Figure 4). This is the attention probability mass assigned by an attention head to a token in the sequence.
Appendix C Path Patching Results on IOI
Figure 6: Full path patching results for the IOI task. Arrows show the three path patching iterations. Color bars show percent logit difference. The lower the number, the more important the head is, as patching it causes the logit difference to drop (i.e., harms performance). The heads that we find most directly affect the logits are the mover heads. Path patching to the mover head query vectors (to see how these heads ‘decide’ where to attend) reveals the inhibition heads (12.3, 13.4, 13.14). Path patching to the value vectors of the inhibition heads (to see what the inhibition heads write into the residual stream), we find the induction and duplicate token heads, which are sensitive to the duplication of the subject token.
In this section we describe the process we use to reverse-engineer the IOI task on GPT2-Medium. We follow the path patching process from Wang et al. (2022) as closely as possible and are able to replicate the same circuit. When choices are made to e.g., path patch to the queries of the mover heads, it is implied that path patching to the values or keys of these
C.1 Mover Heads
@@
Figure 7: Mover head attention patterns showing copying behavior towards the IO token. The negative mover head 19.1 is shown at the bottom with the opposite trend: attending to the subject token and demoting it. A positive correlation means that the more attention a head pays to a token, the more it promotes it as the next word prediction.
C.2 The Negative Mover Head
@@ C.3 Inhibition Heads
Appendix D Path Patching Results on Colored Objects
@@ D.1 Mover Heads
We show examples of mover heads on the Colored Objects task in Figure 11, which shows that the dot product between the token embedding for color words being attended to and the output of the mover head is high, and correlates positively with the attention paid to color words. That is: these heads copy whatever they attend to, and in this setting, they attend to the color words in the test example. Notice that these are the same heads as in IOI (see §C.1).
@@
Figure 12: Showing the effect of blocking the attention from the [end] position of the content gatherer heads to various tokens. Blocking attention to the obj2 token causes the model to degrade to about random chance and is similar to zero ablating the content gatherers. Blocking attention to the token “ color” does not have this effect.
@@
D.2 Content Gatherer Heads
Figure 13: Average attention paid across the entire input sequences by the content gatherer heads from the [end] position. Consistent with our proposed role of these heads, attention is primarily paid to the obj2 position (the object being asked about) and the surrounding context.
In the main paper, we highlight a sample of the tokens that the content gatherer heads attend to from the last token. The average attention paid to all tokens in the prompts from the [end] position is shown in Figure 2. In Figure 12, we show evidence that the content gatherer heads are accounting for approximately all of the signal telling the model which object to predict the color of. Specifically, blocking attention to the obj2 token causes the model to dip to around random performance between the three color options. Zero-ablating the content gatherer heads both give a similar performance drop. Interestingly, blocking attention to the token “ color” does not affect performance negatively at all, but because of the nature of the prompt, there are multiple places that the model can look to gather that the task is to predict a color word, so we can not conclude how important attending to this token actually is for the content gatherer heads.
D.3 What are the Inhibition Heads Doing?
@@
Figure 15: As shown in the main paper, blocking attention in the content gatherer heads from the [end] position to the obj2 token (the object in the question “What color is the obj2?”) reduces performance to randomly selecting between the three color options. This figure shows how blocking this attention redistributes the attention paid by the top five mover heads. When the correct answer is listed as the first or third (index 0 and 2) options listed in the prompt, this reduces attention to the correct answer and increases attention to the wrong answers. This is not the case when the correct answer is the middle option. The model tends to be sensitive to the position of the colors presented, which we discuss in Appendix F.
D.4 What is the Negative Mover Head doing?
@@
Appendix E Duplicate Token and Induction Heads
@@
Appendix F Additional Sources of Error in Colored Objects
@@
F.1 Bias Towards Position in a Sequence
@@
Appendix G Additional Details on Intervention Results
@@
Appendix H Details on Colored Objects Dataset Construction
@@
Appendix I Circuit Overlap in Additional Tasks
@@
I.1 Greater Than Circuit
@@
I.2 World Capitals Circuit:
@@
I.3 Challenges in Quantifying Circuit Overlap
@@
Appendix J Do Other Models also Use Similar Circuits for These Tasks?
@@
Appendix K How was the Colored Objects Task Selected?